# Quantum-thermodynamic enhancements in continuous thermal machines require energetic coherence

### v1

@@ V Thermodynamic impact of coherence
where $J_{\mathrm{out}}^{\mathrm{cl}}$ denotes the output current in the classical-equivalent model. The above ratio measures the relative reduction in the dispersion of the output current in the quantum machine, as compared to the classical one. If $\mathcal{R}>1$, then the classical equivalent provides a more accurate output; $\mathcal{R}=1$ implies that the quantum and classical models are indistinguishable from their dispersion, and $\mathcal{R}<1$ implies that the output in the quantum machine is more accurate, hence providing a quantum-thermodynamic advantage manifested in an enhanced reliability of the machine for same average outputs.
@@
Figure 2: a) Histogram of sampled values of the ratio between the fluctuations of the system and its classical analog $\mathcal{R}$ for different ratios of the bath temperatures (see legend). The values corresponds to an exploration of the following region in the parameters space of the system: $\beta_{\mathrm{c}}=1$, $\epsilon_{2}=5$, $\omega_{\mathrm{d}}\in[0.1,4.9]$, $\gamma_{\mathrm{h}/\mathrm{c}}\in[10^{-5},10^{-2}]$ and $g\in[10^{-5},10^{-2}]$. b) Colour maps of the fluctuation ratio $R$ and c) the TUR ratio $\mathcal{Q}_{\mathrm{he}}$ as a function of the bath interaction strengths and the driving force. The other systems parameters are: $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{h}}=0.1$, $\epsilon_{2}=5$, $\omega_{\mathrm{d}}=2.5$ and $\gamma_{\mathrm{c}}=10^{-3}$.
Our analysis is performed as follows: we first obtain the classical equivalents for the three prototypical machines considered here as explained in Sec. IV. Then we obtain, using the FCS approach, the variance of the output currents for both quantum and classical equivalent models (details are provided in App. B). Finally, we numerically generate a $10^{6}$ number of possible system configurations in the whole parameters space verifying the basic assumptions ensuring the consistency of the Markovian dynamics and conduct a direct comparative analysis.
V.1 Thermodynamic enhancement through energetic coherence
We provide evidence that in thermal machines that show steady-state energetic coherence (as the one induced by a weak external coherent driving) quantum-thermodynamic improvements can be achieved in the form of reduced fluctuations in the output current for given (fixed) power output and efficiency. As a paradigmatic example, we use the three-level amplifier (see Fig. 1a) as introduced in Sec. II.1, to illustrate this point.
@@ V.1 Thermodynamic enhancement through energetic coherence
In order to quantify the impact of (energetic) coherence in the three-level amplifier, we compute the fluctuations ratio $\mathcal{R}$ in Eq. (31) for the output power, $J=\dot{W}$. Exploring the model parameters for fixed external temperatures of the baths, we observe the appearance of a significant amount of configurations with $\mathcal{R}<1$, that increases as the temperature bias powering the machine increases. This is illustrated in Fig. 2a where the distribution of $\mathcal{R}$ values over $10^{6}$ configurations is shown for three different choices of (fixed) environmental temperatures. As can be appreciated, for all configurations we obtain $\mathcal{R}\leq 1$, that is, the three-level amplifier always match or exceed the performance of the corresponding classical equivalent machine for each configuration, hence unveiling a beneficial role of energetic coherence. Improvements reaching an output power variance reduction up to $\mathcal{R}\sim 1/2$ are possible for temperature bias of the order $T_{\mathrm{h}}=10T_{\mathrm{c}}$. We also observe a fat tail in the distribution that ensures the robustness of the enhancements, meaning that many configurations can lead to significant reductions in the output variance. The range of variance reductions shrinks towards higher $\mathcal{R}$ values as the temperature bias is reduced, and tends to disappear close to equilibrium (similar temperatures of the baths) where quantum and classical models perform almost equally. This effect is a manifestation of the nonequilbrium character of the enhancements produced by energetic coherence.
@@
We observe that the area where violations of the TUR, $\mathcal{Q}_{\mathrm{he}}<2$, are verified, is contained within the region $\mathcal{R}<1$ and indeed coincide with the highest improvements in precision as measured by the reduction ratio $\mathcal{R}$. However, as expected, we also find that even in regimes where the TUR is not violated, there exists an improvement in accuracy of the output current due to the presence of energetic coherence. Therefore, using the classical equivalent of the original three level amplifier, we are able to identify regimes of thermodynamic enhancement that cannot be revealed by violations of the TUR.
V.2 Thermodynamic disadvantage caused by degenerate coherence
We now use the same methodology to carry out a similar analysis for the cases of degenerate (non-energetic) coherence, taking as main examples the three-qubit autonomous refrigerator introduced in Sec. IIB and the NIC machine model introduced in Sec. IIC.
@@ V.2 Thermodynamic disadvantage caused by degenerate coherence
Starting with the autonomous refrigerator, we find that the classical equivalent is obtained by replacing the three-body interaction Hamiltonian $V$ allowing the exchange of energy between qubits, by a classical transition producing incoherent jumps between levels |$\ket{101}$ and |$\ket{010}$. The rate of this transition, according to Eq. (26) becomes in this case $\gamma_{\mathrm{uv}}^{\mathrm{cl}}=\gamma_{\mathrm{vu}}^{\mathrm{cl}}={4g^{2}}/\sum_{r}{(2\bar{n}_{r}+1)\gamma_{r}}$, where |$\ket{\mathrm{u}}=\ket{101}$ and |$\ket{\mathrm{v}}=\ket{010}$, and the sum runs over the three baths, $r=\mathrm{h,m,c}$.
@@
Figure 3: Histogram of sampled values of $\mathcal{Q}$ (inset plots) and the ratio between the fluctuations of the system and its classical equivalent $R$ (outset plots). The values corresponds to an exploration of the following region in the parameters space of the system: $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{m}}/\beta_{\mathrm{c}}\in[0,1]$, $\beta_{\mathrm{h}}/\beta_{\mathrm{m}}\in[0,1]$, $\epsilon_{2}=5$, $\epsilon_{1}\in[0.1,4.9]$, $\gamma_{\mathrm{c}/\mathrm{m}/\mathrm{h}}\in[10^{-5},10^{-2}]$ and $g\in[10^{-4},10^{-2}]$.
@@
For assessing the enhancements in the reliability of the refrigerator, the fluctuations ratio $\mathcal{R}$ in Eq. (31) is computed for the cooling power (heat current from the cold bath), $J_{\mathrm{out}}=\dot{Q}_{\mathrm{c}}$. The distributions for the reduction ratio $\mathcal{R}$ and TUR ratio $\mathcal{Q}_{\mathrm{abs}}$ (inset) in this case are shown in Fig. 3 again for $10^{6}$ parameter configurations. As can be observed, in this case we obtain $\mathcal{R}\geq 1$ for all configurations, meaning that a reduction of the fluctuations ratio is not possible. As a consequence, we have to conclude that the three-qubit quantum refrigerator does not perform better than its classical equivalent, despite it has been shown to operate using entanglement [20]. In particular, in the majority of cases, the fluctuations in both systems are comparable, with certain cases where the quantum system exhibits noise levels up to $5\%$ higher than its classical thermodynamic equivalent. Looking at the distribution of $\mathcal{Q}_{\mathrm{abs}}$ values, we also see that the autonomous refrigerator remains unable to break the classical constraint imposed by the TUR in all system configurations (inset plot). Consequently, we conclude that degenerate coherence does not have a significant impact on the autonomous refrigerator performance, which is not able to beat its classical equivalent counterpart in all the parameter regions.

### v2

@@ V Thermodynamic impact of coherence
where $J_{\mathrm{out}}^{\mathrm{cl}}$ denotes the output current in the classical thermodynamic-equivalent model. The above ratio measures the relative reduction in the dispersion of the output current in the quantum machine, as compared to the classical one. If $\mathcal{R}<0$, then the classical equivalent provides a more accurate output than the original quantum one; $\mathcal{R}=0$ implies that the quantum and classical models are indistinguishable from their dispersion, and $\mathcal{R}>0$ implies that the output in the quantum machine is more accurate, hence providing a quantum-thermodynamic advantage manifested in an enhanced reliability of the machine for same average output and efficiency.
Our aim here is to obtain analytical expressions for the variances of the output currents for generic quantum machines and its classical equivalent counterpart by using the full counting statistics formalism, in order to construct the relative difference $\mathcal{R}$ in Eq. (31) and evaluating its sign. This can be performed in principle for both cases of Hamiltonian-induced coherence and noise-induced coherence, following respectively Secs. IV.1 and IV.2. We will first focus on the case of Hamiltonian-induced coherence, since it is in that case that more universal results can be obtained, and then discuss also the case of noise-induced coherence.
As we have seen in Sec. IV a set of rate equations for the evolution of the populations and coherence is be obtained for both the original quantum engine and its classical equivalent counterpart, which for Hamiltonian-induce coherence are given in Sec. IV.1. However, we notice that even assuming a single transition with coherence, the general $N$-level problem is composed by $N-2$ equations of the form:
$\dot{\rho}_{nn}=\sum_{j\neq n}(\gamma_{jn}\rho_{jj}-\gamma_{ni}\rho_{nn})$ (32)
plus 3 more equations for the coherent subspace in the quantum machine, Eqs. (23), or 2 for equations for the classical equivalent, Eqs. (25); with the additional simplification that in steady state conditions, we have $\dot{\rho}_{nn}=0$ $\forall n$. As a consequence, it is clear that in order to obtain analytical results in the generic case, we need to reduce the dimension of the problem. Here, we are able to reduce (and fix) the dimension of the machine by introducing a “coarse grained” mesostate, similar to the one presented in Ref. [80], involving almost all of the incoherent transitions as follows.
Figure 2: Schematic representation of a generic $N$ level thermal machine with Hamiltonian-induced coherence in a subspace composed by two levels (green levels). The transitions between all other levels are produced by thermal baths at possibly different temperatures (simple arrows), and one of its transitions is being monitored (black detector). A reduced version of the machine model is obtained by introducing the coarse-grained state $S$ (purple shaded box) including $N-3$ levels. In the unicycle case the rates of the transitions |$\ket{\rm v}\leftrightarrow\ket{S}$ and |$\ket{\rm u}\leftrightarrow\ket{\rm m}$ are set to zero, as well as the transitions between levels within the coarse-grained state |$\ket{S}$ leading to multiple cycles.
More precisely, we define a “mesostate” |$\ket{S}$ that comprises $N-3$ levels of the thermal machine that are not involved in the coherent interaction (see Fig. 2 for a diagram). We keep outside |$\ket{S}$ only the coherent subspace (levels |$\ket{\rm u}$ and |$\ket{\rm v}$) and an extra level (denoted without loss of generality as |$\ket{\rm m}$ for “monitoring”) connected to it by at least one stochastic transition mediated by a thermal bath (here |$\ket{{\rm v}}\leftrightarrow\ket{{\rm m}}$ without loss of generality), that allow us to properly apply the full counting statistics formalism to uncover the fluctuations in the flux over that transition  [^1] We notice that it is not convenient to apply the FCS formalism to the transitions inside $S$ since, once defined the mesostate, it is not possible anymore to associate an specific energy exchange to any transition from or to $S$..
Being composed by an ensemble of levels, the occupation probability of the mesostate can be written as
$\rho_{SS}:=\sum_{j\in S}\rho_{jj},$ (33)
verifying that $\sum_{i\notin S}\rho_{ii}+\rho_{SS}=1$. Moreover effective rates for the transitions involving the mesostate and any other level $i$ can be defined as
$\begin{split}&\Gamma_{Si}:=\dfrac{1}{p_{S}}\sum_{j\in S}\gamma_{ji}\rho_{jj},~{}~{};~{}~{}\Gamma_{iS}:=\sum_{j\in S}\gamma_{ij},\end{split}$ (34)
where only the rates of transitions going “out” of the mesostate $\Gamma_{Si}$ to the other levels depend on the steady state populations. Then by introducing the mesostate $S$, the $N-2$ set of equations (32) can be reduced to the following 2 equations
$\begin{split}\dot{\rho}_{\rm mm}&=\sum_{i\notin S}\left(\gamma_{i{\rm m}}\rho_{ii}-\gamma_{{\rm m}i}\rho_{\rm mm}\right)+\Gamma_{S{\rm m}}\rho_{SS}-\Gamma_{{\rm m}S}\rho_{\rm mm},\\ \dot{\rho}_{SS}&=\sum_{i\notin S}\left(\Gamma_{iS}\rho_{ii}-\Gamma_{Si}\rho_{SS}\right).\end{split}$ (35)
to be combined with Eqs. (23) for the coherent levels in the quantum machine model, or with Eqs. (25) for the corresponding classical equivalent machine. In any case these extra equations remain unchanged by the substitution of the remaining levels by the mesostate |$\ket{S}$.
Notably, with the above procedure we have fixed the dimension of the general problem to a 4-states thermal machine with coherence between two of these states. That means applying FCS with a matrix of dimension 5 in the quantum case and of dimension 4 for the classical thermodynamic-equivalent machine (see Appendix B). Here it is also important to notice that, by allowing transitions between |$\ket{S}$ and all the other (three) states of the machine (as well as transitions among the levels within $S$), this general procedure applies to thermal machines that can be both unicycle or multicyclic, where the unicycle case is recovered by setting to zero the rates of the transition |$\ket{{\rm v}}\leftrightarrow\ket{S}$ and |$\ket{\rm u}\leftrightarrow\ket{\rm m}$.
The next step is to obtain analytical expressions for the variance of the current in the monitored transition |$\ket{\rm v}\leftrightarrow\ket{\rm m}$. This is done for both the quantum machine and its classical counterpart. Here it is important to mention that, in order to asses the fluctuations of the relevant steady-state currents (including coherent work in the transition |$\ket{\rm u}\leftrightarrow\ket{\rm v}$) from FCS in a single transition of the machine, we have to impose some specific conditions on the transition structure of the machine. In particular, we assume that every transition in the machine is coupled to a single thermal reservoir and that quanta are conserved in all interactions with the reservoirs in the weak-coupling limit, which can be considered as mild conditions. A more restrictive condition is that all thermal reservoirs must contribute the same number of quanta in all the cycles present in the original machine (including those closed by the coherent transitions). This condition (trivially) includes uniclycle machines, but also a broad class of multicycle structures that we dub ”symmetric multiclyce” machines. A relevant example of those structures are obtained in multipartite scenarios like the three-qubit fridge introduced in Sec. II.2, for which all possible cycles involve the exchange of one quantum with each reservoir. This is also the case for many other engines and refrigerators, such as e.g. continuous versions of the SWAP engine [82], fridges made up of bosonic modes [7, 11] or using optomechanical-like couplings [83], as well as quantum dot devices [56, 84], just to mention a few.
In all these situations, the variances of the currents are proportional to each other, ${\rm Var}[\dot{Q}_{r}]=\Delta\epsilon_{r}^{2}{\rm Var}[\dot{N}]$ for all reservoirs $r$ with ${\rm Var}[\dot{N}]:={\rm Var}[\dot{N}_{l}]~{}\forall l$, and also ${\rm Var}[{\dot{W}}]=\omega_{d}^{2}{\rm Var}[\dot{N}]$ for the power output (see also App. B). Therefore, the relative difference in the fluctuations in Eq. (31) can be conveniently rewritten as:
$\mathcal{R}:=\dfrac{\mathrm{Var}[\dot{N}^{\mathrm{cl}}]-\mathrm{Var}[\dot{N}]}{\mathrm{Var}[\dot{N}]},$ (36)
where ${\rm Var}[\dot{N}^{\rm cl}]$ denotes the corresponding variance of the current in the classical equivalent model. The explicit expressions for ${\rm Var}[\dot{N}]$, ${\rm Var}[\dot{N}^{\rm cl}]$ are provided in Appendix D for a generic quantum machine with $N$ levels and $1$ coherent transition. In the case of more than $1$ (independent) coherent transitions, they can be handled by using the same techniques independently for every coherent subspace, all of which are keep outside the coarse-grained state $S$ (hence augmenting the effective dimension of the system as a counterpart).
Once the general expressions for the variances are obtained, the sign of $\mathcal{R}$ can be evaluated to observe its universal character. We resume the result of this analysis enunciating the following theorem, whose explicit demonstration can be found in Appendix E:
@@
Figure 3: a) Histogram of sampled values of the ratio between the fluctuations of the system and its classical analog $\mathcal{R}$ for different ratios of the bath temperatures (see legend). The values corresponds to an exploration of the following region in the parameters space of the system: $\beta_{\mathrm{c}}=1$, $\epsilon_{2}=5$, $\omega_{\mathrm{d}}\in[0.1,4.9]$, $\gamma_{\mathrm{h}/\mathrm{c}}\in[10^{-5},10^{-2}]$ and $g\in[10^{-5},10^{-2}]$. b) Colour maps of the fluctuation ratio $R$ and c) the TUR ratio $\mathcal{Q}_{\mathrm{he}}$ as a function of the bath interaction strengths and the driving force. The other systems parameters are: $\beta_{c}=1$, $\beta_{\mathrm{h}}=0.1/\beta_{c}$, $\epsilon_{2}=5$, $\omega_{\mathrm{d}}=2.5$ and $\gamma_{\mathrm{c}}=10^{-3}$ (energies are given in $k_{\rm B}T_{\rm c}=1$ units).
Theorem 1
Under the presence of Hamiltonian-induced coherence, quantum thermal machines in the weak coupling limit outperform the precision of their classical thermodynamic-equivalent counterparts (as defined as in Sec. IV) in steady-state conditions. In particular it is verified that the relative difference in the fluctuations is always strictly positive $\mathcal{R}>0$ when the systems are in out-of-equilibrium conditions and becomes zero in equilibrium.
The above theorem means that out of equilibrium, the quantum thermal machine is always more precise than its classical thermodynamic equivalent. Since the classical thermodynamic equivalent will have by construction equal average output currents and efficiency than the quantum machine, that implies that Hamiltonian-induced coherence invariably leads to quantum-thermodynamic advantages. In order words, there is no classical model that using the same resources can reproduce the power, efficiency and reliability of the quantum thermal machine.
Theorem 1 is a main result of our fluctuation analysis that applies to generic unicyclic and symmetric multicycle quantum thermal machines operating in nonequilibrium stationary states (NESS) in the weak coupling regime and under weak driving or perturbations, so that their evolution is describable by the Lindlbad master equation, Eq. (2). Examples of such thermal machines are ubiquitous [1], comprising few level masers and lasers under weak driving [47, 12, 85, 86], superconducting devices [87, 88, 33], or many-body engines [89, 32], as well as a number of autonomous engines, refrigerators and quantum clock models [56, 90, 58, 57, 91, 92] composed by various units that interact weakly among them.
The impact of coherence in the case of noise-induced coherence can be also analyzed using a similar method than in the Hamiltonian case. However in this case more complications arise. The set of rate equations describing the evolution of the system are given in Appendix C [Eqs. (70) and (71)]. In that case, a mesostate similar to the one above might be introduced, but it can only include all the machine levels that are not directly connected to the coherent subspace spanned by states ${\rm u}$ and ${\rm v}$. All other state transitions will be now affected by the presence of the coherent transition [c.f. Eq. (29)] and the allowed parameter regimes need to ensure their positivity. Moreover, the intrinsic multiclycic structure no longer ensures that the variances of the steady-state currents in all the transitions are equal as before, but the results may depend on the chosen level to perform the FCS monitoring.
As a consequence, contrary to the Hamiltonian-induced case, the impact of noise-induced coherence turns out to be not universal and depends on the specific model, the specific parameters of the machine, and the specific transition in which FCS in performed, leading to the following:
Theorem 2
Under the presence of noise-induced coherence, quantum thermal machines in the weak coupling limit can either outperform or be outperformed by their classical thermodynamic-equivalent counterparts (as defined as in Sec. IV) in steady-state conditions. In particular the relative difference in the fluctuations $\mathcal{R}$ can be either positive or negative depending on the specific parameters of the machine.
The proof of this theorem is provided in the next section and follows from constructing and analyzing a (counter)example, the NIC machine introduced in Sec. II.3 where regimes with either $\mathcal{R}>0$ and $\mathcal{R}<0$ are founded for different parameter regimes. The importance of Theorem 2 relies in that it puts firmer criteria for the classification of quantum-thermodynamic advantages in noise-induced machines that the ones previously considered in the literature, such as the consideration of the dephased model [16, 17, 18, 19], and going beyond the absence of TUR violations, as it has been checked in some noise-induced models, see e.g. Ref. [68].
In the following section we illustrate our results for the three prototypical machines introduced before, showing the advantages produced by Hamiltonian-induced coherence in the first two models (both for energetic and degenerate cohere) and determining the range of parameters in which the third one (the NIC engine) leads to advantages and disadvantages with respect to its classical equivalent counterpart, suggesting the careful revision of previous claims for quantum-advantages in this type of machines.
VI Illustrative examples
We now illustrate our results for the three paradigmatic examples of quantum machines discussed in Sec. II constructing explicitly their classical equivalent machines and assessing the role of the coherence in the fluctuations of the currents. We numerically generate a $10^{6}$ number of possible system configurations in the whole parameter space, that verify the basic assumptions that ensure the consistency of the Markovian dynamics, and perform a direct comparative analysis between quantum and classical models.
VI.1 Three level amplifier
@@
In order to quantify the impact of (energetic) coherence in the three-level amplifier, we compute the fluctuations ratio $\mathcal{R}$ in Eq. (31) for the output power, $J=\dot{W}$. Exploring the model parameters for fixed external temperatures of the baths, we observe the appearance of a significant amount of configurations with $\mathcal{R}>0$, that increases as the temperature bias powering the machine increases. This is illustrated in Fig. 3a where the distribution of $\mathcal{R}$ values over $10^{6}$ configurations is shown for three different choices of (fixed) environmental temperatures. As can be appreciated, for all configurations we obtain $\mathcal{R}\geq 0$, that is, the three-level amplifier always match or exceed the performance of the corresponding classical equivalent machine for each configuration, hence unveiling a beneficial role of energetic coherence. Improvements reaching an output power variance reduction up to $\mathcal{R}\sim 1$ are possible for temperature bias of the order $T_{\mathrm{h}}=10T_{\mathrm{c}}$. We also observe a fat tail in the distribution that ensures the robustness of the enhancements, meaning that many configurations can lead to significant reductions in the output variance. The range of variance reductions shrinks towards higher $\mathcal{R}$ values as the temperature bias is reduced, and tends to disappear close to equilibrium (similar temperatures of the baths) where quantum and classical models perform almost equally. This effect is a manifestation of the nonequilbrium character of the enhancements produced by energetic coherence.
@@
We observe that the area where violations of the TUR, $\mathcal{Q}_{\mathrm{he}}<2$, are verified, is contained within the region $\mathcal{R}>0$ and indeed coincide with the highest improvements in precision as measured by the reduction ratio $\mathcal{R}$. However, as expected, we also find that even in regimes where the TUR is not violated, there exists an improvement in accuracy of the output current due to the presence of energetic coherence. Therefore, using the classical equivalent of the original three level amplifier, we are able to identify regimes of thermodynamic enhancement that cannot be revealed by violations of the TUR.
VI.2 Three-qubit autonomous refrigerator
@@
For the autonomous refrigerator model, we find that the classical equivalent is obtained by replacing the three-body interaction Hamiltonian $V$ allowing the exchange of energy between qubits, by a classical transition producing incoherent jumps between levels |$\ket{101}$ and |$\ket{010}$. The rate of this transition, according to Eq. (26) becomes in this case $\gamma_{\mathrm{uv}}^{\mathrm{cl}}=\gamma_{\mathrm{vu}}^{\mathrm{cl}}={4g^{2}}/\sum_{r}{(2\bar{n}_{r}+1)\gamma_{r}}$, where |$\ket{\mathrm{u}}=\ket{101}$ and |$\ket{\mathrm{v}}=\ket{010}$, and the sum runs over the three baths, $r=\mathrm{h,m,c}$.
@@
Figure 4: Histogram of sampled values of $\mathcal{Q}$ (inset plots) and the ratio between the fluctuations of the system and its classical equivalent $R$ (outset plots). The values corresponds to an exploration of the following region in the parameters space of the system: $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{m}}/\beta_{\mathrm{c}}\in[0,1]$, $\beta_{\mathrm{h}}/\beta_{\mathrm{m}}\in[0,1]$, $\epsilon_{2}=5$, $\epsilon_{1}\in[0.1,4.9]$, $\gamma_{\mathrm{c}/\mathrm{m}/\mathrm{h}}\in[10^{-5},10^{-2}]$ and $g\in[10^{-4},10^{-2}]$ (energies are given in $k_{\rm B}T_{\rm c}=1$ units).
@@
For assessing the enhancements in the reliability of the refrigerator, the relative fluctuations $\mathcal{R}$ in Eq. (31) is computed for the cooling power (heat current from the cold bath), $J_{\mathrm{out}}=\dot{Q}_{\mathrm{c}}$. The distributions for the reduction ratio $\mathcal{R}$ and TUR ratio $\mathcal{Q}_{\mathrm{abs}}$ (inset) in this case are shown in Fig. 4 again for $10^{6}$ parameter configurations. As can be observed, in this case we obtain $\mathcal{R}\geq 0$ for all configurations, meaning that a reduction of the fluctuations ratio is always possible. As a consequence, we have to conclude that the three-qubit quantum refrigerator performs better than its classical equivalent, despite it has been shown to operate using entanglement [20]. In particular, in the majority of cases, the fluctuations in both systems are comparable, with certain cases where the quantum system exhibits noise levels up to $1.5\%$ lower than its classical thermodynamic equivalent. Looking at the distribution of $\mathcal{Q}_{\mathrm{abs}}$ values, we also see that the autonomous refrigerator remains unable to break the classical constraint imposed by the TUR in all system configurations (inset plot).
VI.3 Noise-induced coherent machine